# An Imitative Reinforcement Learning Framework for Autonomous Dogfight

### v1

@@ 3.2. Deep Learning Methods
Reinforcement learning introduces more autonomous and adaptive methods for dogfight (Wang and Wei, 2022; Fu et al., 2022; Li et al., 2023; Fan et al., 2022; Cao et al., 2023). The RL agent actively interacts with the environment and learns optimal policies through trials and errors. The RL agent adjusts behavior based on received reward signals, gradually improving performance to adapt to the constantly changing environment. Fan et al. (Fan et al., 2022) propose an autonomous maneuver decision method for UCAV based on Asynchronous Advantage Actor-Critic (A3C) (Mnih et al., 2016) algorithm. Relying on a library of seven predefined actions, they train UCAV to gain an advantageous position in aerial combat, and the training efficiency is improved through a multi-threaded asynchronous mechanism. Cao et al. (Cao et al., 2023) integrate RL with game theory, using Double Deep Q-Network (DDQN) (Van Hasselt et al., 2016) algorithm to train a three-degree-of-freedom UCAV to execute seven basic aerial combat maneuvers and maintain a superior position in air combat. Wang et al. (Wang and Wei, 2022) address the issue of multi-UCAV aerial combat by designing reward functions that incorporate prior knowledge. They utilize Multi-Agent Deep Deterministic Policy Gradient (MADDPG) (Lowe et al., 2017) algorithm to overcome the challenges associated with high-dimensional, continuous action spaces that traditional reinforcement learning methods face. This approach enables multiple UCAVs to maintain a superior position in aerial combat. Fu et al. (Fu et al., 2022) use Deep Deterministic Policy Gradient (DDPG) (Lillicrap et al., 2015) algorithm to train a two-degree-of-freedom UCAV to complete tracking and evasion tasks in UCAV dogfight, introducing imitation learning to address the inefficiency of exploring different scenarios during the training phase of reinforcement learning algorithms. Li et al. (Li et al., 2023) build upon Soft Actor-Critic (SAC) (Haarnoja et al., 2018) method, utilizing expert knowledge to prevent SAC from getting stuck in local optima and accelerating the training process. This work trains a three-degree-of-freedom UCAV to perform pursuit and lock decisions in UCAV dogfight, but has not addressed the weapon usage problem. Zhang et al. (ZHANG et al., 2023) propose to expand dogfight decision-making from a single dimension to multiple dimensions, including target pursuit, target lock, and weapon use. This work employs the SAC algorithm for policy learning and establishes a set of meta-policies. However, this work is limited to a two-dimensional environment, and the weapon meta-policies are manually defined.
@@
Although deep learning methods, especially deep RL, have made significant progress in UCAV dogfight, current research still face with several challenges, such as unrealistic simulation environments, oversimplified task design, low algorithmic sample efficiency, and inadequate exploration capabilities. To address these issues, we construct a realistic UCAV dogfight simulation environment based on the Harfang3D sandbox which utilizes a UCAV control mechanism aligned with real operational procedures, and design a comprehensive dogfight task that incorporates multiple decision dimensions. Furthermore, we propose a novel imitative reinforcement learning framework that not only enhances learning efficiency through expert imitation but also improves the policy’s adaptability to dynamic environments through autonomous exploration enabled by reinforcement learning, thereby mastering complex multi-dimensional decision policies in UCAV dogfight tasks.
@@ 4. Problem Formulation
We establish a realistic simulation environment for UCAV dogfight policy learning based on the Harfang3D sandbox (Özbek et al., 2022) and model the ‘pursuit-lock-launch’ tasks in dogfight as an MDP, which is shown in Figure 2.
@@ 4.1. Task Description
In this subsection, we describe the UCAV dogfight task,which is divided into three consecutive stages: target pursuit, target lock, and missile launch. The initial positions of our aircraft and the opponent aircraft are fixed, with a relative distance of $4060$ meters. The opponent aircraft flies forward at a constant speed, while our aircraft is controlled intelligently through a neural network. If our aircraft catches up with the opponent aircraft, locks onto it, launches a missile, and finally shoots down the opponent aircraft, the mission is successful. To ensure that the agent can successfully and accurately complete this task, it needs to master at least three policies of three different stages. Successful target lock requires meeting lock conditions and maintaining them for $5$ seconds, as given below.
@@
where $|\vec{D}|$ is the relative distance between our aircraft and the opponent aircraft, and $q$ is the angle shown in Figure 3(a).
Figure 3. The red aircraft denotes our aircraft, and the blue one denotes the opponent aircraft.
@@ 4.2. State Space
(2) $\vec{D}=(\Delta x,\Delta y,\Delta z)=(x_{u},y_{u},z_{u})-(x_{e},y_{e},z_{e}).$
@@
The state space is comprised of 14 variables: the Cartesian coordinate differences between our aircraft and the opponent aircraft $(\Delta x,\Delta y,\Delta z)$, the Euler angle of our aircraft (in the simulation environment, the three-dimensional Euler angles correspond to the pitch angle, yaw angle, and roll angle, and the corresponding angles can be proportionally converted), heading angle of our aircraft, heading angle of the opponent aircraft, pitch angle of the opponent aircraft, roll angle of the opponent aircraft, locking angle, opponent aircraft health, locking status, and missile status.
@@ 4.4. Reward Function
The reward function is a critical component for applying reinforcement learning to real-world problems. In the UCAV dogfight, the operated aircraft pursues the opponent aircraft, locks its aim, and launches a missile within an extremely short time. To accomplish this task, we propose the following reward function with four terms:
(3) $R=R_{distance}+R_{lock}+R_{altitude}+R_{launch}.$
$R_{distance}$ is a reward function designed to ensure that the operated aircraft can catch up with the opponent aircraft. This reward calculates the distance between the operated and opponent aircraft and uses a negative distance value as the reward function defined as
@@
The reward in Eq.(4) encourages the operated aircraft to swiftly catch the opponent aircraft for subsequent target lock.
$R_{lock}$ is the reward that incentivizes the operated aircraft to lock onto the opponent aircraft. It uses the negative locking angle value for a smaller locking angle to provide a higher reward. The reward function is given by
(5) $R_{lock}=-10q.$
@@
$R_{launch}$ is the reward that encourages a successful launch. This reward is given only when the launch succeeds as follows:
(7) $R_{launch}=\left\{\begin{array}[]{lr}-1,&\text{if $B_{f}=1$ and $B_{m}=0$}\\ 100,&\text{if $B_{f}B_{m}=1$ and $B_{l}=1$}\\ -4,&\text{if $B_{f}B_{m}=1$ and $B_{l}=0$}\\ 0,&otherwise\end{array}\right.$
where $B_{f}$, $B_{m}$, and $B_{l}$ are Boolean values set to $1$ (true) when the aircraft launches a missile, has a loaded missile, and locks onto the opponent, respectively. $R_{launch}$ promotes the aircraft to execute the correct launch action, with the highest reward being achieved when the aircraft launches and has a missile after locking onto the opponent and the lowest reward being achieved when the aircraft launches a loaded missile without locking onto the opponent aircraft.
Among the four reward terms, $R_{distance}$, $R_{lock}$, and $R_{launch}$ facilitate the smooth completion of pursuit, lock, and launch, respectively, while $R_{altitude}$ constrains the exploration space of the operated aircraft to achieve effective exploration.
@@ 5. Methodology
Figure 4. The proposed learning framework for dogfight.
The actor–critic framework serves as the foundation for dogfight policy learning. To mitigate the overestimation of the critic, we adopt a dual Q-network architecture ($Q_{\phi_{1}}$ and $Q_{\phi_{2}}$). Inspired by the RL objective shown in Eq.(8), the actor network $\theta_{actor}$ is trained to maximize the smaller Q value between the two Q values.
(8) $L_{RL}=-E_{s\sim D}[min_{i=1,2}Q_{\phi_{i}}(s,\pi(s;\theta_{actor}))],$
where $L_{RL}$ denotes the RL loss for the actor network, $D$ represents the experience pool, $\pi(s;\theta_{actor})$ denotes the policy network with parameter $\theta_{actor}$, and $Q(s,a)$ represents the estimated Q value by the critic network. Minimizing loss $L_{RL}$ to update the parameters of the actor network can help dogfight policy learning.
@@
where $\mathcal{T}^{e}$ denotes the expert trajectory obtained through artificial intelligence (AI) mode in Harfang3D sandbox and $(s^{e},a^{e})$ denotes the expert state–action pairs. The proposed learning objective, as expressed in Eq.(10), supports agents in autonomous exploration and improves their performance for critical actions in tasks (e.g., launch) to approach expert-level capabilities.
@@
where $\alpha$ is a fixed weight to balance the magnitude $L_{RL}$ and $L_{BC}$ , and $\lambda$ is a weight to balance the two loss terms. In this work, we propose to use the loss function $(1-\lambda)L_{RL}+\alpha\lambda L_{BC}$ to update the actor network, so that we can take the advantage of both RL and imitation learning. However, balancing these two learning objectives poses a significant challenge. To address this issue, we develop two mechanisms to set the balancing factor $\lambda$: linear $\lambda$ and adaptive $\lambda$.
Linear $\lambda$: The initial value for $\lambda$ is $1$, and it gradually decreases with the training round $e$ as follows:
(11) $\lambda=1-e\times k,$
where $k$ is a constant, and we set $k$ as $0.001$ in the experiments.
@@
3:  Initialize weight $\lambda$, replay buffer $D$, and expert trajectories $\mathcal{T}^{e}$
@@
8:        Obtain reward $r$, next state $s^{\prime}$, and add this transition to $D$
@@
The proposed dogfight policy learning framework is outlined in Algorithm 1. First, the actor and critic networks are randomly initialized, and the target networks are synchronized. In lines 7–8, samples are collected in the simulated environment and added to the replay buffer $D$ of the first-in-first-out store manner. In lines 9–12, samples are drawn randomly from the replay buffer $D$, and the actor and critic networks are updated. The loss function for the critic network is as follows:
@@
where $Q_{\phi_{i}^{\prime}}$ represents the target Q networks, and $\hat{a}$ denotes the target action at the next state, $\hat{a}\leftarrow\pi(s^{\prime};\theta_{actor}^{\prime})$. We employ the target networks to stabilize the learning process. As indicated in line 13-14 of Algorithm 1, the target networks are updated in a soft way with weight $\tau$.
@@ 6. Experiments
  • How does the proposed learning framework perform compared with state-of-the-art RL and imitation learning approaches? (Section 6.1)
@@
  • How is the generalization ability of the proposed framework? (Section 6.3)
  • How does the proposed learning framework perform using different kinds of expert data? (Section 6.4)
@@ 6.1. Comparative Results
We compare the proposed learning framework with four state-of-the-art RL and imitation learning methods:
@@
  • Twin-delayed deep deterministic policy gradient (TD3) method (Fujimoto et al., 2018), a RL method designed for continuous action spaces, which improves the deep deterministic policy gradient (DDPG) algorithm (Lillicrap et al., 2015) with double Q value functions.
  • Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018), a maximum entropy RL algorithm.
@@
  • Expert actor-based SAC approach (E-SAC) (Li et al., 2023): The E-SAC algorithm is an imitative RL algorithm. By incorporating high-quality samples generated by experts into the training process of the SAC algorithm, the exploration efficiency and learning speed of the agent are significantly enhanced. Additionally, it can dynamically adjust the proportion of expert samples to exploration samples, ensuring that the learned policy can reach the expert level.
Table 1 lists the parameter settings used in the experiments. The actor and critic networks, besides the input and output layers, consist of two hidden layers with the Leaky Relu activation function. Expert trajectories used by our method are collected in AI mode by $11$ successful episodes, and it contains $13122$ state-action pairs. In contrast, E-SAC method utilizes a larger dataset of expert trajectories, which are collected by BC strategy, and it comprises a total of $100000$ state-action pairs. To enhance the quality of the data collected by BC strategy, we conduct a selection process, adopting only those trajectories that are successful and of high quality.
Table 1. Hyperparameters
Hyperparameter Value
$\alpha$ $1e^{5}$
$\tau$ $0.005$
$T$ $6000$
MaxEpisode $600$
Learning rates $1e^{-3}$
Size of replay buffer $D$ $1e^{6}$
Batch size $128$
Sizes of hidden layers $512$, $256$
Training iterations of BC $1.2e^{7}$
Expert buffer size of our method $1.3e^{4}$
Expert buffer size of E-SAC $1e^{5}$
The experiment results are shown in Figure 5, which depict the success rates and the returns respectively. Every $25$ training episode, we evaluate the learned policies and calculate the average success rate for $20$ episodes. Each curve in Figure 5 represents the average values of four experiments for the same method, with the shaded area indicating the $95\%$ confidence interval.
@@
Figure 5. Comparative results of the proposed approach and the baselines in the dogfight task.
@@
In the highly realistic simulation environment constructed in this study, the agent has mastered two policies for shooting down opponent aircraft: (1) directly ramming the opponent aircraft and (2) hitting the opponent aircraft with a missile. Among them, using missiles to shoot down opponent aircraft resulted in higher reward values, which is the correct decision-making we expect the agent to learn. Figure 5 shows the success rate of shooting down opponent aircraft using either of the two policies.
@@
As shown in Figure 5, various methods employed can converge to high returns in the environment and task we constructed, demonstrating their applicability to reinforcement learning methods. Among them, the adaptive weighting method proposed in this study achieved the highest convergent return, followed by the E-SAC method. Further comparison of the success rate curves between our method and the TD3 method reveals that the imitation of objectives can improve learning efficiency and success rates. This effect is significant when it is adopted in the form of adaptive weights. Moreover, compared to other methods, especially the E-SAC method, which also utilizes expert knowledge, our method not only achieves a higher success rate but also exhibits superior stability. This superiority is primarily attributed to the imitated objectives, which directly optimize the policy network and enable rapid improvement of the policy with small samples in a short time, effectively restricting the agent’s exploration space. This enables the agent to explore along the correct path and gradually master key decision-making skills. On the other hand, the E-SAC method indirectly optimizes the policy and Q-value networks by incorporating expert knowledge into the buffer. Although this approach also utilizes expert knowledge, it requires more expert input and results in lower sample efficiency. Our method minimizes the reliance on extensive expert knowledge while ensuring high stability alongside improving success rates by directly optimizing the policy network. As the imitation learning method, BC, learns in an offline way, so it does not have a learning curve, and therefore we report the convergent performance of BC in the following table.
@@
Table 2 lists the success rates and returns of the best models among the four runs, with data formatted as “mean ± standard deviation”, reflecting the average results and deviation for $50$ evaluation episodes. The table lists both the “success rate” achieved by shooting down opponent aircraft by any means and the “hit success rate” achieved by launching missiles to shoot down opponent aircraft. The code of the proposed approach, the learned models, and recorded videos of the learned policies are available at https://github.com/zrc0622/HIRL4UCAV.
Table 2. Success rates and returns for the best model during learning
Algorithm Success Rate Hit success rate Return
Ours (adaptive) $\bf{100.0\%\pm 0.0\%}$ $\bf{100.0\%\pm 0.0\%}$ $\bf{-680.8\pm 6.7}$
Ours (linear) $\bf{100.0\%\pm 0.0\%}$ $\bf{100.0\%\pm 0.0\%}$ $-958.9\pm 13.8$
TD3 $0.0\%\pm 0.0\%$ $0.0\%\pm 0.0\%$ $-4707.2\pm 0.0$
E-SAC $\bf{100.0\%\pm 0.0\%}$ $\bf{100.0\%\pm 0.0\%}$ $-1431.2\pm 0.2$
SAC $\bf{100.0\%\pm 0.0\%}$ $0.0\%\pm 0.0\%$ $-2985.7\pm 6.2$
BC $62.8\%\pm 1.0\%$ $62.8\%\pm 1.0\%$ $-12228.3\pm 880.2$

### v2

@@ 3.3. Reinforcement Learning
Reinforcement learning introduces more autonomous and adaptive methods for dogfight (wang2022research, ; fu2022uav, ; li2023manoeuvre, ; fan2022air, ; cao2023autonomous, ). The RL agent actively interacts with the environment and learns optimal policies through trial and error. The RL agent adjusts behavior based on received reward signals, gradually improving performance to adapt to the constantly changing environment. Fan et al. (fan2022air, ) propose an autonomous maneuver decision method for UCAV based on the Asynchronous Advantage Actor-Critic (A3C) (mnih2016asynchronous, ) algorithm. Relying on a library of seven predefined actions, they train UCAV to gain an advantageous position in aerial combat, and the training efficiency is improved through a multi-threaded asynchronous mechanism. Cao et al. (cao2023autonomous, ) integrate RL with game theory, using Double Deep Q-Network (DDQN) (van2016deep, ) algorithm to train a three-degree-of-freedom UCAV to execute seven basic aerial combat maneuvers and maintain a superior position in air combat. Wang et al. (wang2022research, ) address the issue of multi-UCAV aerial combat by designing reward functions that incorporate prior knowledge. They utilize the Multi-Agent Deep Deterministic Policy Gradient (MADDPG) (lowe2017multi, ) algorithm to overcome the challenges associated with high-dimensional, continuous action spaces that traditional reinforcement learning methods face. This approach enable multiple UCAVs to maintain advantageous positions in aerial combat. Fu et al. (fu2022uav, ) use Deep Deterministic Policy Gradient (DDPG) (lillicrap2015continuous, ) algorithm to train a two-degree-of-freedom UCAV to complete tracking and evasion tasks in UCAV dogfight, introducing imitation learning to address the inefficiency of exploring different scenarios during the training phase of reinforcement learning algorithms. Li et al. (li2023manoeuvre, ) build upon the Soft Actor-Critic (SAC) (haarnoja2018soft, ) method, utilizing expert knowledge to prevent SAC from getting stuck in local optima and accelerating the training process. This work trains a three-degree-of-freedom UCAV to perform pursuit and lock decisions in UCAV dogfight, but has not addressed the weapon usage problem. Zhang et al. (zhang2023multi, ) propose to expand dogfight decision-making from a single dimension to multiple dimensions, including target pursuit, target lock, and weapon use. This work employs the SAC algorithm for policy learning and establishes a set of meta-policies. However, this work is limited to a two-dimensional environment, and the weapon meta-policies are manually defined.
3.4. Combining Reinforcement Learning and Imitation Learning
To the best of our knowledge, this work is the first to investigate UCAV control policy learning via combining RL and imitation learning. In other domains, similar techniques have been tried. For example, in the offline RL domain, Fujimoto et al. (fujimoto2021minimalist, ) propose to combine the TD3 (fujimoto2018addressing, ) method and the BC method. In contrast to this offline learning method, the proposed approach can leverage online exploration to further enhance policy learning. In the quantitative trading domain, previous researchers combine RL with imitation learning to solve the challenging portfolio management, market making, and order execution problems (imm, ; macmic, ; niu2022metatrader, ).
@@ 3.4. Combining Reinforcement Learning and Imitation Learning
Although deep learning methods, especially deep RL, have made significant progress in UCAV dogfights, current research still faces several challenges, such as unrealistic simulation environments, oversimplified task design, low algorithmic sample efficiency, and inadequate exploration capabilities. To address these issues, we construct a realistic UCAV dogfight simulation environment based on the Harfang3D sandbox, which utilizes a UCAV control mechanism aligned with real operational procedures, and design a comprehensive dogfight task that incorporates multiple decision dimensions. Furthermore, we propose a novel imitative reinforcement learning framework that enhances learning efficiency through expert imitation. Beyond that, the proposed approach can improve the policy’s adaptability to dynamic environments through autonomous exploration, and therefore mastering complex multi-dimensional decision policies in UCAV dogfight tasks.
@@ 4. Problem Formulation
We establish a realistic simulation environment for UCAV dogfight policy learning based on the Harfang3D sandbox (ozbek2022harfang3d, ) and model the ‘pursuit-lock-launch’ tasks in a dogfight as an MDP, which is shown in Figure 2.
@@ 4.1. Task Description
In this subsection, we describe the UCAV dogfight task, which is divided into three consecutive stages: target pursuit, target lock, and missile launch. The initial position of the opponent aircraft is fixed, while the initial position of our aircraft is random, with a relative distance greater than the lock range between them. Both our aircraft and the opponent aircraft are initialized in the air. The opponent aircraft have three modes: flying forward with a constant speed, serpentine maneuvering, and circling maneuvering. In contrast, our aircraft is controlled intelligently through a neural network. If our aircraft catches up with the opponent aircraft, locks onto it, launches a missile, and finally shoots down the opponent aircraft, the mission is successful. To ensure that the agent can successfully and accurately complete this task, it needs to master at least three policies of three different stages. Successful target lock requires meeting lock conditions and maintaining them for $5$ seconds, as given below.
@@
where $|\vec{D}|$ represents the relative distance between our aircraft and the opponent aircraft, and $q$ is the angle shown in Figure 3(a).
Figure 3. An illustration of dogfight rewards. The blue aircraft denotes our aircraft, and the red one denotes the opponent aircraft.
@@ 4.2. State Space
(2) $\vec{D}=(\Delta x,\Delta y,\Delta z)=(x_{e},y_{e},z_{e})-(x_{u},y_{u},z_{u}),$
@@
The state space is comprised of 13 variables: the Cartesian coordinate differences between our aircraft and the opponent aircraft $(\Delta x,\Delta y,\Delta z)$, the Euler angle of our aircraft, locking angle, locking status, missile status, the Euler angle of opponent aircraft, and opponent aircraft health.
@@ 4.4. Reward Function
The reward function is a critical component for applying reinforcement learning to real-world problems. In the UCAV dogfight, the operated aircraft pursues the opponent aircraft, locks its aim, and launches a missile within an extremely short time. To accomplish this task, we propose the following reward function with five terms:
(3) $R=R_{distance}+R_{lock}+R_{success}+R_{altitude}+R_{launch}.$
The five kinds of rewards are labeled in Figure 3 near the corresponding determining variables of these rewards.
$R_{distance}$ is a reward function designed to ensure that the operated aircraft can catch up with the opponent aircraft. This reward calculates the distance between the operated and opponent aircraft and uses a negative distance value as the reward function defined as:
@@
The reward in Eq.(4) encourages the operated aircraft to swiftly close the distance with the opponent aircraft for subsequent target lock.
$R_{lock}$ is the reward that incentivizes the operated aircraft to lock onto the opponent aircraft. It uses the negative locking angle value, where a smaller locking angle provides a higher reward. The reward function is given by:
(5) $R_{lock}=-10q,$
where $q$ is the locking angle between the velocity vector of our aircraft $\vec{V}_{u}$ and the relative distance vector $\vec{D}$.
$R_{success}$ is the final reward for achieving success in air combat. This reward is given only when the missile successfully hits the opponent, as follows:
(6) $R_{success}=\left\{\begin{array}[]{lr}800,&\text{if $B_{s}=1$}\\ 0,&otherwise\end{array}\right.$
where $B_{s}$ is a Boolean value set to $1$ (true) when the missile launched by our aircraft hits the opponent. $R_{success}$ promotes the aircraft to execute the correct launch action, only providing the final reward only when the opponent is shot down by a missile, with no reward given if the opponent is shot down by ramming.
@@
where $z_{u}$ is the altitude of our aircraft as shown in Figure 3.
$R_{launch}$ is the penalty for missile launches. Given that only one missile is available, this penalty encourages the operated aircraft to minimize unnecessary launches. The function is given by
(8) $R_{launch}=\left\{\begin{array}[]{lr}-6,&\text{if $B_{f}=1$}\\ 0,&otherwise\end{array}\right.$
where $B_{f}$ is a Boolean value set to $1$ (true) when our aircraft launches a missile.
Among the five reward terms, $R_{distance}$, $R_{lock}$, and $R_{success}$ facilitate the smooth completion of pursuit, lock, and launch, respectively, while $R_{altitude}$ and $R_{launch}$ constrains the exploration space of the operated aircraft to achieve effective exploration.
@@ 5. Methodology
Figure 4. The proposed learning framework for dogfights.
The actor–critic framework serves as the foundation for dogfight policy learning. Inspired by the TD3 algorithm (fujimoto2018addressing, ), to mitigate the overestimation of the critic, we adopt a double Q-network architecture ($Q_{\phi_{1}}$ and $Q_{\phi_{2}}$) (van2016deep, ). Inspired by the RL objective shown in Eq.(9), the actor network $\theta_{actor}$ is trained to maximize the smaller Q value between the two Q values.
(9) $L_{RL}=-E_{s\sim B}[min_{i=1,2}Q_{\phi_{i}}(s,\pi(s;\theta_{actor}))],$
where $L_{RL}$ denotes the RL loss for the actor network, $B$ represents the replay buffer, $\pi(s;\theta_{actor})$ denotes the policy network with parameter $\theta_{actor}$, and $Q(s,a)$ represents the estimated Q value by the critic network. Minimizing loss $L_{RL}$ to update the parameters of the actor network can achieve dogfight policy learning. The detailed procedure of neural network optimization includes random sampling a minibatch from the replay buffer $B$, and then optimizing $\theta_{actor}$ with stochastic gradient descent. This is standard in the deep learning domain, and it can be regarded as optimizing the expectation in Equation (9). For the rest loss functions, the same procedure is adopted.
@@
where $\mathcal{T}^{e}$ denotes the expert trajectory obtained through artificial intelligence (AI) mode in Harfang3D sandbox and $(s^{e},a^{e})$ denotes the expert state–action pairs. The proposed approach assumes that an effective expert dataset $\mathcal{T}^{e}$ is provided. Without expert data, the proposed approach is downgraded to conventional RL methods.
The proposed learning objective, as expressed in Eq.(11), supports agents in autonomous exploration and improves their performance for critical actions in tasks (e.g., launch) to approach expert-level capabilities.
@@
where $\alpha$ is a fixed weight to balance the magnitude $L_{RL}$ and $L_{BC}$, which is key to learning performance , and $\lambda$ is a weight to balance the two loss terms. In this work, we propose to use the loss function $(1-\lambda)L_{RL}+\alpha\lambda L_{BC}$ to update the actor network, so that we can take advantage of both RL and imitation learning. However, balancing these two learning objectives poses a significant challenge. To address this issue, we develop two mechanisms to set $\lambda$: linear $\lambda$ and adaptive $\lambda$.
Linear $\lambda$: The initial value for $\lambda$ is $0.5$, and it gradually decreases with the training round $e$ as follows:
(12) $\lambda=0.5-e\times k,$
where $k$ is a constant, and we set $k$ as $0.0002$ in the experiments.
@@
3:  Initialize weight $\lambda$, replay buffer $B$, and expert trajectories $\mathcal{T}^{e}$
@@
8:        Obtain reward $r$, next state $s^{\prime}$, and add this transition to $B$
@@
The proposed dogfight policy learning framework is outlined in Algorithm 1. First, the actor and critic networks are randomly initialized, and the target networks are synchronized. In lines 7–8, samples are collected in the simulated environment and added to the replay buffer $B$ of the first-in-first-out store manner. In lines 9–12, samples are drawn randomly from the replay buffer $B$, and the actor and critic networks are updated. The loss function for the critic network is as follows:
@@
where $Q_{\phi_{i}^{\prime}}$ represents the target Q networks, and $\hat{a}$ denotes the target action at the next state, $\hat{a}\leftarrow\pi(s^{\prime};\theta_{actor}^{\prime})$. We employ the target networks to stabilize the learning process. As indicated in Lines 13-14 of Algorithm 1, the target networks are updated in a soft way with weight $\tau$.
@@ 6. Experiments
  • How does the proposed learning framework perform compared with state-of-the-art RL and imitation learning approaches? Can the proposed learning framework work against different types of opponent aircraft? (Section 6.1)
@@
  • Is the proposed framework robust under different conditions? (Section 6.3)
@@ 6.1. Comparative Results
We compare the proposed learning framework with five state-of-the-art RL and imitation learning methods:
@@
  • Twin-delayed deep deterministic policy gradient (TD3) method (fujimoto2018addressing, ), a RL method designed for continuous action spaces, which improves the deep deterministic policy gradient (DDPG) algorithm (lillicrap2015continuous, ) with double Q value functions.
  • Soft Actor-Critic (SAC) algorithm (haarnoja2018soft, ), a maximum entropy RL algorithm.
@@
  • Expert actor-based SAC approach (E-SAC) (li2023manoeuvre, ): The E-SAC algorithm is an imitative RL algorithm. By incorporating high-quality samples generated by experts into the training process of the SAC algorithm, the exploration efficiency and learning speed of the agent are significantly enhanced. Additionally, it can dynamically adjust the proportion of expert samples to exploration samples, ensuring that the learned policy can reach the expert level.
  • Distributional soft actor-critic (DSAC-v2) (duan2023dsac, ): An improved actor-critic approach with distributional value function to enhance the value estimation accuracy.
Table 1 and Table 2 list the parameter settings used in the experiments. Most hyper-parameters are set as the same values for all the methods, as shown in Table 1, and we tune the replay buffer size for each method to achieve their best performance. As for the neural network structure, the actor and critic networks, besides the input and output layers, consist of two hidden layers with the Relu activation function. The expert datasets for all the baseline methods are collected in AI mode with $20$ successful episodes, where the AI mode operates based on a series of predefined rules, such as directing the aircraft towards the opponent and launching missiles when a lock-on is achieved. The expert dataset contains nearly $43e$ 4 state-action pairs, provided at https://github.com/zrc0622/HIRL4UCAV. The episode terminates upon hitting the opponent aircraft down, so the lengths of successful trajectories vary.
Table 1. Common hyper-parameters among all the methods.
Hyper-parameters Value Hyper-parameters Value Hyper-parameters Value
Critic lr 0.001 Actor lr 0.001 Target update interval 3
Batch size 128 Optimizer Adam Training iterations of BC 8e6
$\tau$ 0.005 Discount factor 0.99 Sizes of hidden layers $256$, $512$
$T_{\text{Straight}}$ 1500 $T_{\text{Serpentine}}$ 1500 $T_{\text{Circling}}$ 1900
Table 2. Different hyper-parameters for each method.
Hyper-parameters Ours TD3 SAC E-SAC DSAC-v2
$\alpha$ 10000 / / / /
Size of replay buffer $B$ 1e5 1e5 2e5 2e5 2e5
In the experiments, the opponent behaviors have three modes: straight line, serpentine maneuvering, and circling maneuvering. In the straight mode, the opponent aircraft flies forward at a constant speed. In the serpentine mode, the rudder angle of the opponent repeatedly changes between positive and negative, while the circling maneuvering opponent performs large-angle horizontal turns by adjusting the ailerons and elevators. The experiment results for these three opponent modes are correspondingly shown in the three columns in Figure 5, where the first row demonstrates the return curves, and the second row demonstrates the success rate curves. For every $25$ training episode, we evaluate the learned policies and calculate the average success rate for $50$ episodes. Each curve in Figure 5 represents the average values of four runs for the same method, with the shaded area indicating the $95\%$ confidence interval. The experiments are conducted on the server equipped with Nvidia 2080Ti GPUs, and each run of our approach in Figure 5 takes no more than 18 hours.
@@
Figure 5. Comparative results of the proposed approach and the baselines against various opponents.
@@
The constructed simulator in this study is with high fidelity, since we have designed a realistic action space and task setting. Beyond that, we consider the different properties of various aircraft, and apply real physics models of the aircraft to the simulated engine. In this realistic simulator, the trained agent has mastered two policies for shooting down the opponent aircraft: (1) directly ramming the opponent aircraft and (2) hitting the opponent aircraft with a missile. Among them, using missiles to shoot down opponent aircraft resulted in higher reward values, which is the correct decision-making we expect the agent to learn. The second row of Figure 5 shows the success rate of shooting down opponent aircraft by launching missiles.
@@
As shown in Figure 5, the adaptive weighting method proposed in this study has achieved the highest convergent return and convergent success rate across the three opponent modes. The comparison of the success rate curves between our method and the TD3 method reveals that the imitation of objectives can improve learning efficiency and success rates [^1] The success rate curves for TD3 in subfigure (a) and (c) are invisible, since TD3 can hardly learn any successful policy and the success rates are always zero.. This effect is significant when it is adopted in the form of adaptive weights. Moreover, compared to other methods, especially the E-SAC method, which also utilizes expert knowledge, our method not only achieves a higher success rate but also exhibits superior stability. This superiority is primarily attributed to the imitated objectives, which directly optimize the policy network and enable rapid improvement of the policy with small samples in a short time, effectively restricting the agent’s exploration space. This enables the agent to explore along the correct path and gradually master key decision-making skills. On the other hand, the E-SAC method indirectly optimizes the policy and Q-value networks by incorporating expert knowledge into the buffer. Although this approach also utilizes expert knowledge, it requires more expert input and results in lower sample efficiency. Our method minimizes the reliance on extensive expert knowledge while ensuring high stability alongside improving success rates by directly optimizing the policy network. As the imitation learning method, BC, learns in an offline way, so it does not have a learning curve, and therefore we report the convergent performance of BC in Table 3.
Table 3. Best-performing results averaged over four runs.
Algorithm Straight line opponent Serpentine opponent Circling opponent
Hit success rate Return Hit success rate Return Hit success rate Return
Ours(adaptive) $\bf{100.0\%\pm 0.0\%}$ $\bf{72.3\pm 26.0}$ $\bf{100.0\%\pm 0.0\%}$ $\bf{62.5\pm 16.9}$ $\bf{96.5\%\pm 1.7\%}$ $\bf{-1187.3\pm 290.3}$
Ours(linear) $\bf{100.0\%\pm 0.0\%}$ $36.6\pm 15.8$ $\bf{100.0\%\pm 0.0\%}$ $36.4\pm 33.0$ $96.0\%\pm 2.4\%$ $-1312.5\pm 316.9$
TD3 $0.0\%\pm 0.0\%$ $-1033.3\pm 67.3$ $25.0\%\pm 43.3\%$ $-1020.9\pm 320.5$ $0.0\%\pm 0.0\%$ $-2233.3\pm 211.2$
SAC $99.0\%\pm 1.7\%$ $30.5\pm 86.3$ $98.5\%\pm 2.6\%$ $-270.6\pm 378.1$ $76.0\%\pm 16.3\%$ $-2029.7\pm 1350.1$
E-SAC $\bf{100.0\%\pm 0.0\%}$ $-248.9\pm 356.0$ $72.5\%\pm 42.0\%$ $-601.8\pm 431.8$ $73.5\%\pm 33.4\%$ $-1598.8\pm 661.1$
DSAC-v2 $0.0\%\pm 0.0\%$ $-791.1\pm 69.6$ $0.0\%\pm 0.0\%$ $-825.5\pm 34.3$ $0.0\%\pm 0.0\%$ $-1197.8\pm 137.9$
BC $\bf{100.0\%\pm 0.0\%}$ $-23.3\pm 15.4$ $73.0\%\pm 9.2\%$ $-1102.7\pm 379.9$ $65.0\%\pm 9.1\%$ $-3055.3\pm 555.9$
Expert $100.0\%\pm 0.0\%$ $9.6\pm 16.6$ $100.0\%\pm 0.0\%$ $4.3\pm 13.7$ $100.0\%\pm 0.0\%$ $-931.3\pm 154.5$
@@
We select the best-performing results from four runs and calculate their average. The results are presented in Table 3. The data in the table is displayed in the format ”mean ± standard deviation” for clarity. The code of the proposed approach, the learned models, and recorded videos of the learned policies are available at https://github.com/zrc0622/HIRL4UCAV.